OpenStudy (anonymous):

Suppose y varies directly with x, and y = 8 when x = –6. What direct variation equation relates x and y? What is the value of y when x = –2? A. y = –0.75x; 1.50 B. y = –1.33x; 2.67 C. y = 1.33x; –2.67 D. y = 0.13x; –0.25

OpenStudy (goformit100):

▬Sir/Ma'am, A Warm Welcome To Open Study. Sir/Ma'am please show your Working for the question you have posted.▬

OpenStudy (anonymous):

i have no idea how to do this\

OpenStudy (zzr0ck3r):

\(y=k*x\) for some \(k\in \mathbb{Z}\) Now we can use the point given to figure out what \(k\) is. We use the point \((-6,8)\). \(8=-6*k\implies k = -\frac{4}{3}\) So we have \(y = -\frac{4}{3}x\) Now plug in \(x=-2\) to get the last part.

OpenStudy (anonymous):

how would u simplify 2 2/3

OpenStudy (anonymous):